# Intensity-Efficient Allocations

### v5

@@ Abstract
This paper proposes a refinement of Pareto-efficient allocations for situations where, in addition to having ordinal preferences, agents also have ordinal intensities: they can make comparisons such as “I prefer $a$ to $b$ more than I prefer $c$ to $d$, without necessarily being able to quantify them. A rank-based criterion for interpersonal comparisons of such ordinal intensities is introduced for this new analytical environment. Building on this, an allocation is defined to be intensity-efficient if it is Pareto-efficient with respect to the agents’ preferences and also such that when another Pareto-efficient allocation assigns the same pairs of items to the same pairs of agents but in a “flipped” way, the former allocation assigns the commonly preferred item in every such pair to the agent who prefers it more. Conditions are established under which such Pareto-refining allocations exist. The potential usefulness of this theory in matching problems is illustrated with a quadratic-time extension of the Random Priority (RP) algorithm that returns an allocation which intensity-dominates RP’s Pareto-efficient one.
@@ 1 Introduction
Introducing and analyzing such a principle, in turn, requires clarity on what is meant by preference intensities. Most of economic theory has traditionally taken the existence of intensity comparisons as synonymous to the existence of a cardinal utility function, and has treated the way in which utility differences derived from such a function are ranked as the determinant of how the individual’s intensity comparisons are ordered. Building on cardinal utility functions over goods that are derived from expected-utility preferences over lotteries over goods (von Neumann and Morgenstern, 1947), and assuming that utility differences are interpersonally comparable, the classical utilitarian summation criterion formalized in Harsanyi (1955) is the benchmark distributional principle that refines the set of Pareto-efficient allocations in this analytical environment (Roemer, 1996; Moulin, 2004).
@@
Despite the undeniable tractability and widespread application of this and other utilitarian criteria in various domains of economic analysis, it is well-understood that the specific cardinal utility functions emerging from expected-utility preferences over lotteries over goods confound an individual’s intensity comparisons with that person’s attitudes toward risk (Arrow, 1951, 1963; Luce and Raiffa, 1957; Baccelli and Mongin, 2016; Sen, 2017; Dietrich, 2025; Samuelson, 2025). More importantly, when cardinal utility differences —even with riskless axiomatic primitives—are used to define an agent’s intensity comparisons, the agent in question is portrayed as having a superhumanly high degree of precision in their capacity to quantify those intensities (Samuelson, 1938; Basu, 1982). This stands in contrast to some contemporary views in the field of cognitive psychology and decision-making, according to which “the perceptual system might be like a pan in balance, which responds by tipping to the left or right, depending on which of two items is heavier, but provides no read-out of the absolute weight of either item.” (Vlaev et al., 2011, p. 551). To the extent that comparing options is easier and more reliable than assigning values to them, as the above survey and several studies cited therein suggest, it is reasonable to hypothesize that comparing preference-intensity differences in a similar fashion is also cognitively less demanding than assigning numerical values to them, hence possibly a more accurate reflection of the individual’s true intensity comparisons.
@@
Motivated by this tension between the intuitive appeal of including intensity information in the allocation process and what might be regarded as weaknesses of the standard cardinal-welfarist approach toward doing so, this study introduces and analyses a novel third domain for social-choice and matching-theoretic analysis which provides a middle ground for further progress to be made. Specifically, instead of cardinal-welfarist or Arrowian profiles that, respectively consist of interpersonally comparable cardinal utility functions—possibly not arising from expected-utility preferences (Nebel, 2024; Weymark, 2005; Harvey, 1999)—or preference orderings, this domain comprises profiles of generally ordinal preference intensity orderings, i.e. collections of quaternary relations that reflect the agents’ preference intensity comparisons. This analytical choice assumes (and this assumption is validated in Section 5, Proposition 3) that, similar to the way in which preferences are elicited, e.g. in matching problems, information about agents’ intensities is obtainable by asking them to respond to simple comparison questions such as “Do you prefer $Apples$ to $Berries$ more than you prefer $Pears$ to GrapesGrapes?”, and that these responses are internally consistent in a way that we make precise in Section 2.
@@
This new and general framework where both preference and preference-intensity information is available to the social planner naturally raises the question of how such information might be used—as per the goals that we set out above—to arrive at some normatively appealing refinement of Pareto-efficiency that would also reflect differences in the agents’ preference intensities and mitigate the multiplicity of Pareto-efficient allocations. Similar to utilitarianism and related cardinal-welfarist criteria, such a refinement requires some kind of interpersonal comparisons. Unlike those notions, however, where such comparisons are based on agents’ utilities (Baccelli, 2023; Echenique et al., 2023; d’Aspremont, 2018; Fleurbaey and Hammond, 2004), in our framework they must rely instead on the information contained in the above ordinal intensity rankings.
@@
To make progress against this analytical challenge, and focusing on the assignment problem with finitely many indivisible goods, we submit that such comparisons can be made by contrasting the rank-order position of pairs of alternatives $(a,b)$ in the different agents’ intensity orderings, whenever such contrasting is meaningful. In particular, when both agents $i$ and $j$ prefer $a$ to $b$ but the pair $(a,b)$ lies higher in $i$’s intensity ranking than in $j$’s, and these rankings have the same length, then we interpret this as suggesting that $i$ prefers it more. In Section 3 we formalize and defend this Interpersonal Comparisons of Ordinal Intensities new postulate/informational basis for social choice and welfare. We also clarify that, unlike utilitarian environments where interpersonal comparisons are driven by agents’ absolute intensities that are captured by the size of the respective utility differences, this postulate instead is rooted in comparing agents’ relative intensities in a scoring-like fashion that is analogous in this environment to the well-known Borda-count (Borda, 1781) or relative-utilitarian (Dhillon and Mertens, 1999) normalizations in ordinal- and cardinal-preference environments, respectively (Baccelli, 2023; d’Aspremont, 2018).
@@
In Section 4 we build on this comparability postulate and introduce a refinement of Pareto efficiency that incorporates such interpersonal comparisons of ordinal intensities, with the goal of arriving at allocations that are more distributively just. Specifically, we define an allocation $x$ to be intensity-efficient if it is Pareto-efficient and also such that, whenever another Pareto-efficient allocation $y$ assigns the same pairs of objects to the same pairs of agents but in a “flipped” way, i.e. when $(x_{i},x_{j})=(y_{j},y_{i})=(a,b)$ for agents $i$, $j$ and alternatives $a$, $b$, then $x$ assigns the commonly preferred alternative in every such pair to the agent who prefers it more. Assuming strict preferences and intensities, we first show that an intensity-efficient allocation exists under every profile, without further restrictions, when there are three agents and alternatives. Yet we also show that with four or more existence is not guaranteed because the underlying dominance relation may be cyclic without further assumptions. For the case of $n$ agents and items we then establish a sufficient condition which requires profiles to be (effectively) homeo-monotonic and whose key part could be summarized informally thus: when a group of $k$ agents have identical preferences over their top $k$ items, and agents can be weakly ordered in how much they prefer any two consecutively-ranked items in this subset, then this ordering cannot be violated in how much they prefer any non-consecutively ranked items. Furthermore, although this condition is not necessary for existence, we prove that a weaker version of it is indeed satisfied by every profile that has an intensity-efficient allocation.
@@
Intensity-efficiency appears to be the first intensity-inclusive refinement of Pareto-efficiency that is operational in an environment where neither the agents’ utilities are required to be intra- or inter-personally comparable nor monetary transfers between agents are assumed to be feasible. Yet an interesting by-product of the existence investigation is the realization that, in those analytical environments where all agents’ intensity comparisons emerge from cardinal utility functions that are assumed to be interpersonally comparable, classical-utilitarian and intensity-efficient allocations do not generally coincide, although they do so under an intuitive additional condition that we introduce. The potential discrepancy between these two intensity-incorporating welfare concepts that may arise in such cardinal environments is itself of potential usefulness in applications, because either notion can be used as a secondary selection criterion in those situations where the other one predicts multiple allocations.
@@
In Section 5, finally, we turn to the problem of designing a matching mechanism that assigns an intensity profile to an allocation that improves upon Pareto-efficiency in the direction of intensity-efficiency. We first show that once agents’ ordinal preferences have been elicited, their ordinal intensities that induce those preferences can be elicited efficiently (specifically, in quadratic time) by leveraging the monotonic “lateral consistency” relation that exists between preferences and intensities in the class we study. Then, building on the matching-theoretic literature with one-sided preferences, we study an algorithm that extends the fundamental Random Priority (RP) one and produces—in quadratic time in the number of items/agents—an allocation that intensity-dominates the Pareto-efficient allocation that is found by RP. While RP makes the elicitation of agents’ preferences incentive-compatible, we show that this does not carry over to the elicitation of agents’ intensities under the proposed extension of the algorithm. We connect our analysis to related recent work in this literature that started with Abdulkadiroğlu, Che, and Yasuda (2011) and which focuses on the trade-off between utilitarian welfare and incentive-compatibility that is relevant in a designer’s choice between ordinal and cardinal matching mechanisms.
1.1 Related Literature
@@
This paper’s contribution lies at the intersection of welfare economics, social choice and matching theory. Similar to existing cardinal-welfarist refinements of Pareto-efficiency such as those based on classical or relative utilitarianism (Harsanyi, 1955; Dhillon and Mertens, 1999), Nash social welfare (Nash, 1950; Luce and Raiffa, 1957; Kaneko and Nakamura, 1979; Roemer, 1996; Moulin, 2004), Pareto-efficient egalitarian-equivalence (Pazner and Schmeidler, 1978) or relative egalitarianism (Kalai and Smorodinsky, 1975; Imai, 1983; Sprumont, 2013), the proposed refinement is inclusive of information that goes beyond agents’ preferences toward also reflecting their intensities in the allocation process. Unlike these refinements, the hereby proposed one is the first that is operational in the general class of ordinal preference-intensity profiles, and builds on a novel interpersonal-comparability criterion that is distinct from interpersonal comparisons of utility differences, which the above notions rely on. Also unlike these refinements, however, intensity-efficient allocations are introduced and studied here directly in the context of the assignment problem, rather than as the application of some social welfare function that is generally defined on the relevant domain. The latter problem is important and deserves separate treatment.
Intensity-efficient allocations and their domain of application are also distinct from those dictated by the Borda scoring rule (Borda, 1781; Young, 1974; Maskin, 2025). Although this rule is defined by a particular normalization of agents’ ordinal utility functions over alternatives, in practice it is often used with a cardinal interpretation due to the fact that ordering the utility differences induced by this normalization defines the special class of intensity relations where every agent prefers two alternatives that are ranked $k$ places apart exactly as much as she prefers any two other alternatives that are ranked $k$ places apart. The class of ordinal intensity relations that we consider does not impose this restriction. Yet at the same time, for the main results of this paper we assume a strictness condition on agents’ intensities which rules out such “linear” intensity profiles.
@@
In addition to Abdulkadiroğlu, Che, and Yasuda (2011), finally, several studies in matching theory—both with one- and two-sided preferences—have taken agents’ interpersonally comparable von Neumann-Morgenstern cardinal utilities (Budish, 2011; Lee and Yariv, 2018; Ortoleva, Safonov, and Yariv, 2023) or quasi-linear preferences in willingness to pay (Che, Gale, and Kim, 2013) as inputs in the allocation process toward maximizing utilitarian welfare. Our study shares the same goal as those above, namely to improve upon the distributive-justice properties of Pareto-efficiency in the final matching, yet pursues it in the unexplored environment where only ordinal intensity comparisons are generally available to the designer. While this environment is less tractable for the modelling analyst, it is simpler for the economic agent and, as such, we believe worthy of further investigation.
2 Notation, Preliminaries and Measurement-Theoretic Foundations
@@
By $A:=\{a_{1},\ldots,a_{n}\}$ and $V:=\{v_{1},\ldots,v_{n}\}$ we denote, respectively, the finite sets of choice alternatives and agents. To simplify the assignment problem under study we assume that these sets have equal sizes. When subscripts are unnecessary we also write $a,b\in A$. Similarly, when no confusion arises we will refer to agent $v_{l}$ by their subscript, $l$. The preference intensity relation of agent $l\leq n$ on $A$ is denoted by $\dot{\succsim}_{l}$, with $\dot{\succ}_{l}$ and $\dot{\sim}_{l}$ its asymmetric and symmetric parts. In line with extensive-measurement theory (Krantz et al., 1971; Pfanzagl, 1971; Roberts, 1979), these are quaternary relations on $A$ or, equivalently, binary relations on $A\times A$. The statement $(a,b)\,\dot{\succsim}_{l}\,(c,d)$ will be interpreted as “$a$ is preferred to $b$ at least as much as $c$ is to dd” when the first option in each pair is the (possibly weakly) preferred one at that pair, and as “$b$ is preferred to $a$ no more than $d$ is preferred to cc” when the converse is true (naturally, $(a,b)\dot{\succsim}_{l}(c,d)$ will also hold when $a$ is preferred to $b$ and $d$ is preferred to $c$). When no ambiguity arises, we will interpret $(a,b)\,\dot{\succsim}_{l}\,(c,d)$ more succinctly as suggesting that the intensity difference between $a$ and $b$ is weakly larger than that between $c$ and $d$.
@@
The definitions that follow omit existential quantifiers, yet they should be understood as imposing conditions that apply to all objects in the respective domains.
@@
A relation $\dot{\succsim}_{l}$ has a utility-difference representation if there is a function $u_{l}:A\rightarrow\mathbb{R}$ that satisfies
@@
In an influential article, psychologist S. S. Stevens (1946) distinguished between four scales of measurement that are afforded by some numerical assignment on a set of items: nominal (qualitative/categorical measurement); ordinal (invariance up to a strictly increasing transformation); interval (invariance up to positive affine transformation); and ratio (invariance up to a positive linear transformation). Utility-difference representations were added to this list by psychologist C. H. Coombs (1950) as those that correspond to ordered metric scale measurement. These are more general than neoclassical cardinal utility representations (Lange, 1934; Alt, 1936, 1971; Samuelson, 1938) because $u_{l}$ in the latter case is an interval scale whereas in the former it “falls logically between an interval scale and an ordinal scale” (p. 145). Baccelli (2024) has recently clarified, however, that every ordinal utility representation defines a subset of utility-difference comparisons whose order is indeed preserved by all ordinal transformations of the original representation. In our environment, whenever $u_{l}$ in (1) is not cardinal it has the complex uniqueness property of additive representations on finite sets.
@@
$(\mathbf{2_{\mathit{D}}})$ [Cancellation]. $(b_{i},c_{i})\,\dot{\succsim}_{l}\,(b_{\pi(i)},c_{\sigma(i)})$ for $i<t$, $i>0$, implies $(b_{\pi(0)},c_{\sigma(0)})\,\dot{\succsim}_{l}\,(b_{0},c_{0})$ for all sequences $b_{0},b_{1},\ldots,b_{t-1},c_{0},\ldots,c_{t-1}\in A$ and all permutations $\pi,\sigma$ of $\{0,\ldots,t-1\}$, where $t>0$.
@@
A relation $\dot{\succsim}_{l}$ is representable by a preference intensity function if there is a mapping $s_{l}:A\times A\rightarrow\mathbb{R}$, unique up to a strictly increasing transformation, such that
$(a,b)\,\dot{\succsim}_{l}\,(c,d)$ $\Longleftrightarrow$ $s_{l}(a,b)\geq s_{l}(c,d)$ (4)
$(a,b)\,\dot{\succsim}_{l}\,(b,a)$ $\Longleftrightarrow$ $s_{l}(a,b)\geq s_{l}(e,e)\geq s_{l}(b,a)$ (5)
$\min\{s_{l}(a,b),s_{l}(b,c)\}\geq s_{l}(e,e)$ $\Longrightarrow$ $s_{l}(a,c)\geq\max\{s_{l}(a,b),s_{l}(b,c)\},$ (6)
with a strict inequality on the left hand side of (6) implying a strict inequality on the right.
In words, $s_{l}$ represents the agent’s intensity weak ordering [(4)] and the preferences induced by it [(5)], further ensuring that these preferences are themselves weakly ordered and intensities are monotonically increasing in the preference ordering [(6)]. Such a representation is possible if and only if $\dot{\succsim}_{l}$ is a weak order on $A\times A$ that additionally satisfies $(\mathbf{3_{\mathit{D}}})$ and the following condition, which states that $a$ is preferred to $b$ if and only if the intensity difference between $a$ and $c$ is higher than that between $b$ and $c$.
@@
The (pseudo-)cardinal utility-difference model is nested by the ordinal model in the special case where “lateral consistency” [(6)] is strengthened to “additivity” [(7)]:

### v6

# Intensity-Efficient and Intensity-Positional Allocations

@@ Abstract
This paper studies the allocation problem in environments where, in addition to having ordinal preferences, agents also have ordinal intensities: they can say “I prefer $a$ to $b$ more than I prefer $c$ to $d$ without necessarily being able to claim how much more. A rank-based criterion for interpersonal comparisons of ordinal intensities is introduced for this analytical environment. Building on it, the concepts of intensity-dominated, intensity-efficient and intensity-positional allocations are defined.Two allocations are related by intensity dominance if they assign the same two items to the same two agents but in a “flipped” way, and one allocation assigns the commonly preferred item in any such pair to the agent who prefers it more. An allocation is intensity-efficient if it is Pareto-efficient and not intensity-dominated. Distinctly, an allocation is intensity-positional if it maximizes the agents’ total intensity-rank scores. The latter extend standard Borda scores to environments where consecutively-ranked items do not necessarily feature identical intensity differences. The paper analyses intensity-efficient and intensity-positional allocations, and clarifies how they differ from those based on the utilitarian or Borda rules.
@@ 1 Introduction
Introducing and analyzing such a principle, in turn, requires clarity on what is meant by preference intensities. Most of economic theory has traditionally taken the existence of intensity comparisons as synonymous to the existence of a cardinal utility function, and has treated the way in which utility differences derived from such a function are ranked as the determinant of how the individual’s intensity comparisons are ordered. Building on cardinal utility functions over goods that are derived from expected-utility preferences over lotteries over goods (von Neumann and Morgenstern, 1947), and assuming that utility differences are interpersonally comparable, the classical utilitarian summation criterion, formalized in Harsanyi (1955), is the benchmark distributional principle that refines the set of Pareto-efficient allocations in this analytical environment (Moulin, 1988, 2004).
@@
Despite the undeniable tractability and widespread application of this and other utilitarian criteria in various domains of economic analysis, it is well-understood that the specific cardinal utility functions emerging from expected-utility preferences over lotteries over goods confound an individual’s intensity comparisons with that person’s attitudes toward risk (Arrow, 1951, 1963; Luce and Raiffa, 1957; Baccelli and Mongin, 2016; Sen, 2017; Alaoui and Penta, 2025; Samuelson, 2025). More importantly, when cardinal utility differences —even with riskless axiomatic primitives—are used to define an agent’s intensity comparisons, the agent in question is portrayed as having a superhumanly high degree of precision in their capacity to quantify those intensities (Samuelson, 1938; Basu, 1982). This stands in contrast to some contemporary views in the field of cognitive psychology and decision-making, according to which “the perceptual system might be like a pan in balance, which responds by tipping to the left or right, depending on which of two items is heavier, but provides no read-out of the absolute weight of either item.” (Vlaev et al., 2011, p. 551). To the extent that comparing options is easier and more reliable than assigning values to them, it is reasonable to hypothesize that comparing preference-intensity differences in a similar fashion is also cognitively less demanding than assigning numerical values to them, hence possibly a more accurate reflection of the individual’s true intensity comparisons.
@@
Motivated by this tension between the intuitive appeal of including intensity information in the allocation process and what might be regarded as weaknesses of the standard cardinal-welfarist approach toward doing so, this study introduces and analyses a novel third domain for social-choice and matching-theoretic analysis which provides a middle ground for further progress to be made. Specifically, instead of cardinal-welfarist or Arrowian profiles that, respectively consist of interpersonally comparable cardinal utility functions—possibly not arising from expected-utility preferences (Nebel, 2024; Weymark, 2005; Harvey, 1999)—or preference orderings, this domain comprises profiles of generally ordinal preference intensity orderings, i.e. collections of quaternary relations that reflect the agents’ preference intensity comparisons. This analytical choice assumes that, similar to the way in which preferences are elicited, e.g. in matching problems, information about agents’ intensities is obtainable by asking them to respond to simple comparison questions such as “Do you prefer $Apples$ to $Berries$ more than you prefer $Pears$ to GrapesGrapes?”, and that these responses are internally consistent in a way that we make precise in Section 2.
@@
This framework where both preference and—generally only ordinal—intensity information is available to the social planner naturally raises the question of how it might be used, as per the goals that we set out above, to arrive at some normatively appealing refinement of Pareto-efficiency that would also reflect differences in the agents’ intensities and mitigate the multiplicity of Pareto-efficient allocations. Similar to utilitarian and related cardinal-welfarist criteria, such a refinement requires some kind of interpersonal comparisons. Unlike those notions, however, where such comparisons are based on agents’ utilities (Fleurbaey and Hammond, 2004; d’Aspremont, 2018; Echenique et al., 2023; Baccelli, 2023), in our framework they must rely instead on the information contained in the above ordinal intensity rankings.
@@
To make progress against this analytical challenge, and focusing on the assignment problem with finitely many indivisible goods, we submit that such comparisons can be made by contrasting the rank-order position of pairs of alternatives $(a,b)$ in the different agents’ intensity orderings, whenever such contrasting is meaningful. In particular, when both agents $i$ and $j$ prefer $a$ to $b$ but the pair $(a,b)$ lies higher in $i$’s intensity ranking than in $j$’s, and these rankings have the same length, then we interpret this as suggesting that $i$ prefers it more. In Section 3 we formalize and motivate this interpretational assumption concerning interpersonal comparisons of ordinal intensities in the present analytical environment. We also clarify that, unlike utilitarian environments where interpersonal comparisons are driven by agents’ absolute intensities that are captured by the size of the respective utility differences, this postulate instead is rooted in comparing agents’ relative intensities in a scoring-like fashion that is analogous in this environment to the well-known Borda-count (Borda, 1781; Young, 1974; Mas-Colell et al., 1995; Maskin, 2025) or relative-utilitarian (Dhillon and Mertens, 1999) normalizations in ordinal- and cardinal-preference environments, respectively (d’Aspremont, 2018; Baccelli, 2023).
@@
In Section 4 we build on this comparability postulate and introduce a refinement of Pareto efficiency that incorporates such interpersonal comparisons of ordinal intensities, with the goal of arriving at allocations that are more distributively just. Specifically, we define an allocation $x$ to be intensity-efficient if it is Pareto-efficient and also such that, whenever another Pareto-efficient allocation $y$ assigns the same pairs of objects to the same pairs of agents but in a “flipped” way, i.e. when $(x_{i},x_{j})=(y_{j},y_{i})=(a,b)$ for agents $i$, $j$ and alternatives $a$, $b$, then $x$ assigns the commonly preferred alternative in every such pair to the agent who prefers it more. Assuming strict preferences and intensities, we first show that an intensity-efficient allocation exists under every profile when there are three agents and alternatives. We also show that existence is not guaranteed when there are more agents and items, because the underlying dominance relation may be then be cyclic.
Motivated by this, for the general case of $n$ agents and as many items we establish a sufficient condition for existence, requiring profiles to be (effectively) homeo-monotonic. The key part of this condition could be summarized informally as follows: when a group of $k$ agents have identical preferences over their top $k$ items, and agents can be weakly ordered in how much they prefer any item in this subset over that item’s immediate successor, then this ordering cannot be violated when it comes to comparing the agents’ intensities between any non-consecutively ranked items in this subset. Importantly, although the presence of such groups—and the corresponding restrictions within them—is not necessary for existence, we show that if a profile does give rise to such a group, and if the above-mentioned ordering over its agents is strict, then homeo-monotonicity in that group is indeed necessary for an intensity-efficient allocation to exist.
@@
Intensity-efficiency appears to be the first intensity-inclusive refinement of Pareto-efficiency that is operational in an environment where neither the agents’ utilities are required to be intra- or inter-personally comparable nor monetary transfers between agents are assumed to be feasible. Yet an interesting by-product of the existence investigation is the realization that, in those analytical environments where all agents’ intensity comparisons do emerge from cardinal utility functions that are assumed to be interpersonally comparable, classical-utilitarian and intensity-efficient allocations do not generally coincide. They do so, however, when the said utility profile satisfies an intuitive additional condition that we introduce. This requires any interpersonal dominance that emerges by comparing two agents’ utility differences to be accompanied by an interpersonal dominance in the same direction in the rank-order positions of these utility differences in the respective agents’ intensity orderings.
In Section 5 we build on a stronger interpersonal comparisons assumption to propose and analyse the concept of intensity-positional allocations. Toward defining them, we recall first than an item’s Borda score for some agent is the number of items that are ranked below it by that agent. Although Borda scores give rise to a specific normalized ordinal utility representation of the agent’s preferences, in practice they are often used with a cardinal interpretation (see, for example, Maskin, 2025) due to the fact that ordering the utility differences induced by this normalization defines the special class of intensity relations where the agent prefers two alternatives that are ranked $k$ places apart exactly as much as she prefers any two other alternatives that are ranked $k$ places apart. We will refer to these intensity relations as “linear”. Similar to Borda allocations, which are defined as those that maximize the aggregate Borda score, intensity-positional allocations are defined as the maximisers of the aggregate intensity-rank score. Unlike an item’s Borda score, however, which reveals the position of that item on the agent’s preference ordering, its intensity-rank score corresponds to the highest rank-order position in the agent’s intensity ordering where this item appears as preferred to something else. Hence, it encodes information both about the number of alternatives that are ranked below this item and about the relative intensity differences between it and the items it is preferred to. This information, in turn, is reflected in the prescribed allocation process. In the paper’s focal case of strict intensity profiles we show that intensity-positional allocations refine standard Borda allocations. We further show that, in a larger domain where a like-for-like comparison remains meaningful, intensity-rank and Borda-score aggregation diverge in general but coincide at linear profiles. The intuitive reason is that, unlike intensity-positional aggregation, Borda aggregation treats every profile as it if were linear. Thus, the intensity-positional criterion is a genuine generalization of the standard Borda-count criterion in this extended domain.
Intensity-positional and intensity-efficient allocations are logically distinct. Two advantages of the former relative to the latter are their generic existence and their incorporation of a way to resolve intensity-pertinent welfare trade-offs across agents. A disadvantage is that making intensity-positional allocations operational implicitly invokes stronger interpersonal comparability assumptions.
@@
In Section 6, finally, we turn to the problem of designing a matching mechanism that assigns an intensity profile to an allocation that improves upon the Pareto-efficiency desideratum. We first show that once agents’ ordinal preferences have been elicited, their ordinal intensities that induce those preferences can be elicited in quadratic time by leveraging the monotonic “lateral consistency” relation that exists between preferences and intensities in the class that we focus on. Then, building on the matching-theoretic literature with one-sided preferences, we study an algorithm that extends the fundamental Random Priority (RP) mechanism’s algorithm and produces—in quadratic time in the number of items/agents—an allocation which intensity-dominates the Pareto-efficient one that is found by RP. While RP makes the elicitation of agents’ preferences incentive-compatible, we show that this does not carry over to the elicitation of agents’ intensities under the proposed extension of the algorithm. We connect our analysis to related recent work in this literature that started with Abdulkadiroğlu, Che, and Yasuda (2011) and focuses on the trade-off between utilitarian welfare and incentive-compatibility that is relevant in a designer’s choice between ordinal and cardinal matching mechanisms.
@@
As far as the broader related literature is concerned, this paper’s contribution lies at the intersection of welfare economics, social choice and matching theory, in all three cases building bridges between the existing ordinal- and cardinal-welfarist paradigms. Similar to existing cardinal-welfarist refinements of Pareto-efficiency such as those based on classical or relative utilitarianism (Harsanyi, 1955; Dhillon and Mertens, 1999), Nash social welfare (Nash, 1950; Luce and Raiffa, 1957; Moulin, 2004) and others, the proposed refinement is inclusive of information that goes beyond agents’ preferences toward also reflecting their intensities in the allocation process. Unlike these refinements, the hereby proposed ones are the first that are operational in the general class of ordinal preference-intensity profiles, and build on novel interpersonal-comparability criteria that are distinct from interpersonal comparisons of utility differences which the above notions rely on. Also unlike these refinements, however, intensity-efficient allocations are introduced and studied here directly in the context of the assignment problem, rather than as the application of some social welfare function that is generally defined on the relevant domain. Intensity-positional allocations on the other hand are indeed derived from a social welfare function. The problem of characterizing this function axiomatically is important and deserves separate treatment.
@@
In addition to Abdulkadiroğlu, Che, and Yasuda (2011), finally, several studies in matching theory—both with one- and two-sided preferences—have taken agents’ interpersonally comparable von Neumann-Morgenstern cardinal utilities (Budish, 2011; Lee and Yariv, 2018; Ortoleva, Safonov, and Yariv, 2023) or quasi-linear preferences in willingness to pay (Che, Gale, and Kim, 2013) as inputs into the allocation process toward maximizing utilitarian welfare. Our study shares the same goal as those above, namely to improve upon the distributive-justice properties of Pareto-efficiency in the final matching, yet pursues it in the unexplored environment where only ordinal intensity comparisons are generally available to the designer. While this environment is less tractable for the modelling analyst, it is simpler for the economic agent and, as such, we believe worthy of further investigation.
2 Measurement-Theoretic Foundations
@@
By $A:=\{a_{1},\ldots,a_{n}\}$ and $N:=\{1,\ldots,n\}$ we denote, respectively, the finite sets of choice alternatives and agents. To simplify the assignment problem under study we assume that these sets have equal sizes. When subscripts are unnecessary we also write $a,b\in A$. Similarly, when no confusion arises we will refer to agent $v_{l}$ by their subscript, $l$. The preference intensity relation of agent $l\leq n$ on $A$ is denoted by $\dot{\succsim}_{l}$, with $\dot{\succ}_{l}$ and $\dot{\sim}_{l}$ its asymmetric and symmetric parts. In line with extensive-measurement theory (Krantz et al., 1971; Pfanzagl, 1971; Roberts, 1979), these are quaternary relations on $A$ or, equivalently, binary relations on $A\times A$. The statement $(a,b)\,\dot{\succsim}_{l}\,(c,d)$ will be interpreted as “$a$ is preferred to $b$ at least as much as $c$ is to dd” when the first option in each pair is the (possibly weakly) preferred one at that pair, and as “$b$ is preferred to $a$ no more than $d$ is preferred to cc” when the converse is true (naturally, $(a,b)\dot{\succsim}_{l}(c,d)$ will also hold when $a$ is preferred to $b$ and $d$ is preferred to $c$). When no ambiguity arises, we will interpret $(a,b)\,\dot{\succsim}_{l}\,(c,d)$ more succinctly as suggesting that the intensity difference between $a$ and $b$ is weakly larger than that between $c$ and $d$.
@@
The definitions that follow omit universal quantifiers, yet they should be understood as imposing conditions that apply to all objects in the respective domains.
@@
[…]
@@
In an influential article, psychologist S. S. Stevens (1946) distinguished between four scales of measurement that are afforded by some numerical assignment on a set of items: nominal (qualitative/categorical measurement); ordinal (invariance up to a strictly increasing transformation); interval (invariance up to positive affine transformation); and ratio (invariance up to a positive linear transformation). Utility-difference representations were added to this list by psychologist C. H. Coombs (1950) as those that correspond to ordered metric scale measurement. These are more general than neoclassical cardinal utility representations (Lange, 1934; Alt, 1936, 1971; Samuelson, 1938) because $u_{l}$ in the latter case is an interval scale whereas in the former it “falls logically between an interval scale and an ordinal scale” (p. 145). [^1] More specifically, it has the complex uniqueness property of additive representations on finite sets, articulated in Krantz et al. (1971, Theorem 2, p. 431). Baccelli (2024) has recently clarified, however, that every ordinal utility representation defines a certain subset of utility-difference comparisons whose order is, in fact, preserved by all ordinal transformations of the original representation.
@@
$(\mathbf{2_{\mathit{D}}})$ [Cancellation]. $(b_{i},c_{i})\,\dot{\succsim}_{l}\,(b_{\pi(i)},c_{\sigma(i)})$ for all $i<t$, $i>0$, implies $(b_{\pi(0)},c_{\sigma(0)})\,\dot{\succsim}_{l}\,(b_{0},c_{0})$ for all sequences $b_{0},b_{1},\ldots,b_{t-1},c_{0},\ldots,c_{t-1}\in A$ and all permutations $\pi,\sigma$ of $\{0,\ldots,t-1\}$, where $t>0$.
@@
[…]
$(a,b)\,\dot{\succsim}_{l}\,(c,d)$ $\Longleftrightarrow$ $\widehat{s}_{l}(a,b)\geq\widehat{s}_{l}(c,d)$ (4)
$(a,b)\,\dot{\succsim}_{l}\,(b,a)$ $\Longleftrightarrow$ $\widehat{s}_{l}(a,b)\geq\widehat{s}_{l}(e,e)\geq\widehat{s}_{l}(b,a)$ (5)
$\min\{\widehat{s}_{l}(a,b),\widehat{s}_{l}(b,c)\}\geq\widehat{s}_{l}(e,e)$ $\Longrightarrow$ $\widehat{s}_{l}(a,c)\geq\max\{\widehat{s}_{l}(a,b),\widehat{s}_{l}(b,c)\},$ (6)
with a strict inequality on the left part of (6) implying a strict inequality on the right.
In words, $\widehat{s}_{l}$ represents the agent’s intensity weak ordering [(4)] and the preferences induced by it [(5)], further ensuring that these preferences are themselves weakly ordered and intensities are monotonically increasing in the preference ordering [(6)]. Such a representation is possible if and only if $\dot{\succsim}_{l}$ is a weak order on $A\times A$ that additionally satisfies $(\mathbf{3_{\mathit{D}}})$ and the following condition, which states that $a$ is preferred to $b$ if and only if the intensity difference between $a$ and $c$ is higher than that between $b$ and $c$.
@@
The ordinal model nests the (pseudo-)cardinal utility-difference model in the special case where “lateral consistency” [(6)] is strengthened to “additivity” [(7)]: